# Tracking the Feature Dynamics in LLM Training: A Mechanistic Study

### v2

@@ 4.1 Understanding Feature Formation from a Geometric Perspective
where $\mathbf{\widehat{W}}_{i}=\mathbf{W}^{\text{enc}}_{i,:}$ and $\widehat{b}_{i}=b^{\text{enc}}_{i}-c\cdot\mathbf{W}^{\text{enc}}_{i,:}\cdot\mathbf{b}^{\text{dec}}$. Using this, we formalize the “feature region” for feature $i$ as follows:
The function $f_{i}(\mathbf{x})$ naturally divides the activation space into regions via the ReLU function, where $\mathbf{\widehat{W}}_{i}\cdot\mathbf{x}+\widehat{b}_{i}>0$ defines a region for $\mathbf{x}$. However, this division ignores the impact of activation strength, which is critical for semantic fidelity. As noted by (Bricken et al., 2023), higher activation levels often indicate stronger associations with specific tokens or concepts. In Fig. 2 (left), a 2D toy example shows how activation levels refine these regions, enabling a more precise characterization:
Definition 4.1 (Feature Region with Activation Level).
A region corresponding to feature $i$ with activation level $[L,U)$ is defined as:
@@
To assess semantic fidelity, we assume a threshold $L_{i}$ beyond which a feature is considered semantically meaningful. This is formalized as:
Definition 4.2 (Faithful Feature Region).
A region that faithfully represents feature $i$ is defined as:
$\mathcal{R}_{i}^{\text{faithful}}=\mathcal{R}_{i}^{[L_{i},\infty)},$ (9)
where $L_{i}$ is the threshold activation level required for semantic fidelity.
In Fig. 2 (right), a toy example demonstrates how the faithful regions of three features partition a 2D activation space. This example highlights the SAE’s objective to identify these regions and provides insights into the geometric structure of the activation space.
Feature formation describes how datapoints with similar semantics converge into faithful feature regions in the activation space. To study this phenomenon, we track the evolution of datapoints within the same faithful feature region. Using the SAE at the final checkpoint, we define $\{\mathcal{R}_{i}^{\text{faithful}}[t=T_{\text{final}}]\}_{i=0}^{F-1}$ as the ground truth. Let $\mathcal{D}_{i}$ denote the set of datapoints that activate within $\mathcal{R}_{i}^{\text{faithful}}[T_{\text{final}}]$ at the final checkpoint, i.e.,
$\mathcal{D}_{i}=\{(\mathcal{C},q)\mid F^{(T_{\text{final}})}(\mathcal{C},q;\Theta^{(T_{\text{final}})})\in\mathcal{R}_{i}^{\text{faithful}}[T_{\text{final}}]\},$ (10)
where $\mathcal{C}$ represents the context and $q$ denotes the token position within $\mathcal{C}$. The activation set of these datapoints at training step $t$ is defined as:
@@
Here, $\mathcal{A}_{i}^{t}$ represents the activations at checkpoint $t$ for all datapoints in $\mathcal{D}_{i}$. Thus, studying feature formation involves analyzing the dynamics of $\{\mathcal{A}_{i}^{t}\}_{i=0}^{F-1}$ across training.
Figure 4: Progress Measure of activation space (Left) and feature space (Right) for Pythia-410m-deduped (Top) and Stanford CRFM GPT-2 Small Alias (Bottom) across training steps (log scale). For discussion of the glitch in Stanford CRFM GPT-2 Small Alias, see Appendix I.
4.2 Visualization and Analysis
By tracking $\{\mathcal{A}_{i}^{t}\}_{i=0}^{F-1}$ across training steps $t$, we visualize the process of feature formation. Using UMAP (McInnes et al., 2020), we project the activation sets of various features across checkpoints (see Fig. 3, more detailed below). Datapoints with the top 25 feature activations are considered faithful, highlighting the dynamics of both token-level and concept-level features.
Initialization and Warmup Phase: Token-level feature datapoints are clusters, while concept-level feature datapoints remain scattered. In Fig. 3, Checkpoint 0, the initialized network shows tightly clustered token-level features. This arises because datapoints $(\mathcal{C},q)$, corresponding to identical tokens in similar contexts $\mathcal{C}$, produce highly similar activations when passed through a transformer $F$ with randomized parameters $\Theta_{\text{rand}}$. The similarity is driven by shared tokens and repetitive contexts. In contrast, concept-level features remain scattered, as abstraction and generalization across related tokens do not occur in a randomized transformer. This phase reflects the early stage of feature development, dominated by basic token-level associations and noise.
Emergent Phase: Concept-level features begin clustering. As training progresses (Fig. 3, Checkpoints 5, 10, and 15), concept-level features transition from scattered points to cohesive clusters, reflecting the model’s growing ability to represent concepts. Token-level features remain tightly clustered and stable throughout this phase.
Convergent Phase: Features stabilize into cohesive clusters. By Fig. 3, Checkpoints 60 and 153, both token-level and concept-level features are stable and well-formed, reflecting the model’s fully developed ability to represent tokens and broader semantic concepts.
These visualizations at Fig. 3 across various checkpoints vividly illustrate the evolution of features, highlighting their transition from early random activations to semantically meaningful and well-defined clusters.
4.3 Progress Measure
It remains unclear whether feature formation represents a phase transition or a progressive process. To address this, we propose the Feature Formation Progress Measure, which quantifies the degree to which a feature becomes well-formed during training. This measure compares the similarity within semantic datapoints to a baseline derived from randomly sampled, unrelated datapoints.
@@ 4.3 Progress Measure
where $\mathcal{A}_{\text{random}}$ represents a set of randomly sampled datapoints, and:
@@
To extend the analysis to feature space, we define a corresponding measure:
where:
Here, $\mathcal{F}_{i}^{t}$ represents the feature set of datapoints encoded by the SAE encoder at training step $t$, while $\mathcal{F}_{\text{random}}$ denotes the feature set for randomly sampled datapoints encoded in the same manner. By incorporating the feature space, this measure provides a finer-grained perspective on how datapoints align with features, offering valuable insights into the process of feature formation.
$Sim(\cdot,\cdot)$ can be computed using various metrics, such as cosine similarity or Jaccard similarity. In Fig. 4, we employ cosine similarity to demonstrate the utility of the proposed Progress Measure in capturing the gradual formation of features. Experiments with alternative similarity measures (e.g., Jaccard similarity) yield consistent results, as detailed in Appendix G, validating the robustness of this metric.
Notably, our Progress Measure captures the dynamic evolution of token-level and concept-level features, offering a quantitative view on feature formation. This complements the qualitative insights discussed in Sec. 4.2. Also, this measure further reveals that token-level and concept-level features are not entirely distinct. There is a continuous spectrum from typical token-level to high-level concept-level features, with intermediates (such as weak concept-level features) existing between them, as detailed in Appendix C.
@@ 5 Analysis of Feature Drift
It remains obscure whether feature directions drift throughout training or remain fixed from early on. Our study reveals that feature directions initially exhibit significant drift, continue to drift even after features are considered “formed,” and ultimately stabilize to their final state. A feature is considered formed once it gains semantic meaning, which typically stabilizes its semantic association. However, our analysis shows that directional drift persists during training, reflecting ongoing refinements in the feature geometry before reaching stability.
@@ 5.1 Decoder Vector Evolution
(b) Cosine similarity progression for sampled features across training checkpoints. Each subplot traces how individual features progressively approach their final direction.
Figure 5: Decoder Vector Evolution. Distribution and progression of cosine similarity between decoder vectors at intermediate checkpoints and their final state, illustrating both global alignment trends and individual feature dynamics throughout training. Conducted on Pythia-410m-deduped.
Our analysis examines the evolution of decoder vectors $\{\mathbf{W}_{:,i}^{\text{dec}}\}_{i=0}^{F-1}$, which represent an overcomplete basis for the activation space. These vectors define the geometric representation of each feature and its contribution to reconstructing the activations. Following (Templeton et al., 2024), the direction of a feature $i$ is defined as $\tfrac{\mathbf{W}^{\text{dec}}_{:,i}}{\left\|\mathbf{W}^{\text{dec}}_{:,i}\right\|}$.
To study the evolution of feature directions, we calculate the cosine similarity between decoder vectors at different training checkpoints. This quantifies how closely the directions of features at time $t$ align with their final state at $t=\text{final}$.
Our analysis uses two complementary perspectives. Fig. 5(a) provides a checkpoint-centric view, summarizing the cosine similarity of all features at a given training time relative to their final state. This perspective offers a global snapshot of the model’s progress in refining feature directions, revealing that feature directions in LLMs do not remain static during training. Instead, most features exhibit significant drift and gradually align with their final state as training progresses.
In contrast, Fig. 5(b) adopts a feature-centric view, illustrating how the similarity of a specific feature with its final state evolves across all training checkpoints. This visualization shows that the similarity initially remains relatively stable, then experiences a sudden increase during training, and gradually transitions into a plateau as it approaches the end. These trends align with the three phases we previously identified: Initialization & Warmup, Emergent, and Convergent. Together, these two perspectives complement each other, providing consistent yet distinct insights into the directional changes of features during training.
Figure 6: Feature Trajectories. Trajectories of decoder vectors represent the directional change of features across training checkpoints. “Dark red” indicates features that are considered “formed,” i.e., they have gained semantic meaning and generally remain stable semantic meaning until the final state. “Blue” indicates features that are still unformed or in the initial stage. Conducted on Pythia-410m-deduped.
@@ 5.2 Trajectory Analysis
Building on the analysis of decoder vector evolution, we now examine feature trajectories to gain a deeper understanding of how features evolve across training checkpoints. By distinguishing the phases before and after a feature gains semantic meaning, we can better understand the relationship between directional drift and feature formation. This distinction highlights how drift persists even after feature formation, providing insights into the continuous adjustments in the feature space throughout training.
@@
Let $\mathbf{W}_{:,i}^{\text{dec}}[t]$ denote the decoder vector for feature $i$ at training checkpoint $t$, where $t\in\{1,\dots,T_{\text{final}}\}$. The trajectory of feature $i$ is defined as:
@@
A feature is considered formed once it gains and stabilizes in semantic meaning. This means that trajectories exhibiting (“maintaining”) behavior are always in the formed state, whereas those exhibiting (“shifting”) or (“grouping”) behaviors transition into the formed state later during training and remain in this state until the final checkpoint.
Fig. 6 illustrates that a feature would continue to drift even after gaining semantic meaning and considered formed, unveiling the dynamic relationship between directional drift and feature formation.
@@ 6 Conclusion
In this paper, we conduct a comprehensive mechanistic analysis of feature evolution in LLMs during training: (1) We propose SAE-Track, a novel method for obtaining a continual series of SAEs across training checkpoints, enabling efficient and stable feature tracking. (2) We mechanistically investigate feature formation and develop a progress measure for it, tracking the progression of features from noise to semantic fidelity. (3) We analyze feature drift, showing that feature directions initially exhibit significant adjustments and continue to drift even after features gain semantic meaning, before ultimately stabilizing at the final checkpoint. Our work provides a detailed understanding of how features evolve throughout training and lays the foundation for future studies. A more detailed discussion on its broader impact on this field, as well as potential future research directions, is provided in Appendix J.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Detailed Derivation of the Training-Step Continuity Theorem

### v3

@@ 4.1 Geometric Perspective and Qualitative Illustration of Feature Formation
where $\mathbf{\widehat{W}}_{i}=\mathbf{W}^{\text{enc}}_{i,:}$ and $\widehat{b}_{i}=b^{\text{enc}}_{i}-c\cdot\mathbf{W}^{\text{enc}}_{i,:}\cdot\mathbf{b}^{\text{dec}}$. Using this, we can define a feature region based on activation strength.
The function $f_{i}(\mathbf{x})$ naturally divides the activation space into regions via the ReLU function, where $\mathbf{\widehat{W}}_{i}\cdot\mathbf{x}+\widehat{b}_{i}>0$ defines a region for $\mathbf{x}$. However, this division ignores the impact of activation strength, which is critical for semantic fidelity. As noted by bricken2023towards , higher activation levels often indicate stronger associations with specific tokens or concepts. We formally define such a region as follows:
Definition 4.1 (Feature Region by Activation Strength).
A region corresponding to feature $i$ with activation strengths in the range $[L,U)$ is defined as:
@@
Figure 3: Feature Formation: Geometric Concepts and UMAP Visualization. Top Left: Feature regions defined by varying activation levels. Bottom Left: Distinct feature regions derived from high activations, highlighting the partitioning within the activation space. Right: UMAP visualizations of activation set evolution for various features (Pythia-410m-deduped), from initial randomness (Checkpoint 0) to coherent clusters (Checkpoint 153), distinguishing token-level (circular) and concept-level (diamond) features.
Fig. 3 (Top Left) provides a 2D toy example showing how varying activation levels, i.e., different values for $L$ in Definition 4.1, can define more precise regions. To focus our analysis on semantically meaningful feature activations, we typically consider datapoints that elicit high activation strengths for a given feature $i$. In practice, this is often achieved by identifying the set of input contexts and tokens $(\mathcal{C},q)$ that produce the top $k$ highest activation values for feature $i$ (e.g., $k=25$ in our visualizations and analyses). This set of top-$k$ activating datapoints effectively defines the most salient region of activity for feature $i$. Fig. 3 (Bottom Left) conceptually illustrates how such distinct high-activation regions for different features can partition the activation space, highlighting the SAE’s objective to identify these separable, semantically rich areas.
Feature formation, therefore, describes how datapoints with similar semantics converge into these high-activation regions associated with specific features. To study this phenomenon, we first identify at the final checkpoint ($T_{\text{final}}$) the set $\mathcal{D}_{i}$ comprising the (context, token) pairs $(\mathcal{C},q)$ that yield the top $k$ activations for each feature $i$. The activation set for these specific $(\mathcal{C},q)$ pairs at any given training step $t$ is then defined as:
@@
Here, $\mathcal{A}_{i}^{t}$ represents the activations at checkpoint $t$ for this consistently defined set of $\mathcal{D}_{i}$ datapoints. Thus, studying feature formation involves analyzing the dynamics of $\{\mathcal{A}_{i}^{t}\}_{i=0}^{F-1}$ across training.
To visually inspect the convergence of activation sets $\{\mathcal{A}_{i}^{t}\}_{i=0}^{F-1}$, we use UMAP projections mcinnes2020umapuniformmanifoldapproximation (see Fig. 3, Right). Specifically, these plots qualitatively depict the geometric convergence of feature activations, where initially dispersed concept-level features (diamonds) progressively form cohesive clusters throughout training, contrasting with token-level features (circles) that often exhibit some degree of clustering from early stages (e.g., Checkpoint 0). While offering valuable intuition, these UMAP plots serve primarily as a qualitative illustration, motivating the rigorous quantitative analysis of feature formation presented next.
4.2 Quantitative Analysis with a Progress Measure
To address the question of whether feature formation is a phase transition or a progressive process, and to provide a more objective assessment than visual inspection (Fig. 3(b)), we propose the Feature Formation Progress Measure. This metric quantifies the degree to which a feature becomes well-formed during training by comparing the similarity within its representative datapoints (identified via top-$k$ activations at $T_{\text{final}}$, as per Sec. 4.1) to a baseline derived from randomly sampled, unrelated datapoints.
@@ 4.2 Quantitative Analysis with a Progress Measure
where $\mathcal{A}_{i}^{t}$ is the activation set of the top-$k$ datapoints for feature $i$ at step $t$, $\mathcal{A}_{\text{random}}$ represents a set of randomly sampled datapoints, and:
@@
We also define a corresponding measure $M_{i}^{\text{feature}}(t)$ in feature space using the SAE-encoded features $\mathcal{F}_{i}^{t}$ and $\mathcal{F}_{\text{random}}$, providing a finer-grained perspective (see definition in Sec. D).
Figure 4: Feature Dynamics Across Training Steps. Progress measures for the activation space (Left) and feature space (Right) of Pythia-410m-deduped across training steps (log scale). The top row shows the dynamics of features with manual annotation as examples, including concept-level and token-level ones. The bottom row presents a broader set of randomly sampled features, categorized into token-level (blue), (weak) concept-level (light red), and concept-level (dark red).
The Progress Measure, applied using cosine similarity, offers quantitative insights into the mechanisms of feature formation. Fig. 4 quantitatively demonstrates these dynamics in both activation and feature spaces, revealing several key characteristics of this process. Specifically:
  • Most features undergo predominantly a gradual formation process, rather than an abrupt transition, evidenced by the smooth evolution of their Progress Measure curves in Fig. 4.
  • Fig. 4 verifies the initial existence of token-level features and the progressive learning of abstract concept-level features (mentioned in Sec. 3). As shown in Fig. 4 (bottom row), token-level features (blue curves) typically exhibit high value from the start, contrasting with concept-level features (dark red curves) which generally start lower and increase gradually throughout training.
  • Concept features follow a three-phase development process (Initialization/Warmup, Emergent, Convergent at Sec. 3). This is quantitatively supported by the characteristic three-phase ’low-rise-stable’ trajectory of their Progress Measure curves in Fig. 4 (e.g., bottom-row dark red curves and several top-row examples), with the plateau often being more pronounced in the feature space plots (right column).
  • A potential spectrum exists between purely token-level and abstract concept-level features, where weak concept-level features (light red curves) like morphological features (top-row examples) often display patterns intermediate to the distinct token-level (blue) and concept-level (dark red) ones (further discussed in Appendix F).
@@ 5 Analysis of Feature Drift
(b) Cosine similarity progression for sampled features to their final direction across training checkpoints. Each colored line represents an individual feature.
Figure 5: Decoder Vector Evolution. (a) Distributions show the global trend of feature direction alignment over training. (b) Individual feature trajectories illustrate the phased nature of this alignment. All analyses conducted on Pythia-410m-deduped.
It remains elusive whether feature directions (i.e., crucial geometric components identified by SAEs) undergo significant drift throughout training or become fixed early on. Understanding this dynamic is vital for a comprehensive picture of how feature representations stabilize. Accordingly, this section investigates the evolution of these feature directions, leading to two important conclusions about their dynamics during the training process.
@@ 5.1 Decoder Vector Evolution
We examine the evolution of feature directions by analyzing their normalized decoder vectors $\tfrac{\mathbf{W}^{\text{dec}}_{:,i}}{\left\|\mathbf{W}^{\text{dec}}_{:,i}\right\|}$ (following templeton2024scaling ) and calculating their cosine similarity with their respective final directions at $t=\text{final}$.
Conclusion 3.1: Decoder Vector Dynamics Our analysis of decoder vector evolution reveals that most feature directions undergo significant drift throughout training, challenging assumptions of early stability. This alignment to their final states is also a three distinct phases that similar to the features’ semantic evolution.
The evidence supporting these conclusions is presented in Fig. 5. The widespread and significant nature of feature drift is demonstrated by Fig. 5(a). These distributions of cosine similarities between current and final feature directions, initially broad and skewed towards lower values, markedly shift towards 1 as training progresses, indicating a global, gradual alignment process.
The three-phase pattern of this alignment is evident in Fig. 5(b), which displays the cosine similarity progression for a collection of sampled features. These trajectories collectively illustrate: (1) an Initialization & Warmup phase, often with low or stable similarity; (2) an Emergent phase, marked by a rapid increase in similarity as features orient towards their final directions; and (3) a Convergent phase, where alignment plateaus at high values. As noted, these dynamic phases in directional convergence similar to the semantic evolution stages discussed in Section 3.
@@ 5.2 Trajectory Analysis
To gain a more granular understanding of how individual feature directions evolve continuously, we build upon the preceding analysis by examining their full trajectories across training checkpoints. This requires defining both the trajectory itself and the concept of a “formed” feature.
@@
Let $\mathbf{W}_{:,i}^{\text{dec}}[t]$ denote the decoder vector for feature $i$ at training checkpoint $t$, where $t\in\{1,\dots,T_{\text{final}}\}$. The trajectory of feature $i$ is defined as the sequence of its decoder vectors:
@@
Definition 5.2 (Formed Feature).
A feature is considered formed once it acquires and maintains a stable semantic meaning. Features exhibiting “maintaining” behavior (Section 3) are considered formed throughout their observed duration. Features undergoing “shifting” or “grouping” are deemed formed from the checkpoint at which they attain a stable semantic role that persists until $T_{\text{final}}$.
Our analysis of these feature trajectories (Definition 5.1) reveals a counter-intuitive yet crucial insight into their stabilization dynamics, as summarized in Conclusion 4. Specifically, we find that the semantic stabilization of a feature does not necessarily coincide with the stabilization of its geometric direction.
Conclusion 3.2: Drift of Semantically Stable Features Trajectory analysis reveals a counter-intuitive dynamic: individual feature directions often continue to drift and geometrically reorganize even after their semantic meaning has stabilized (i.e., they are “formed”). This underscores that while a feature’s semantic role may crystallize, its precise directional embedding undergoes further refinement until the overall feature ensemble converges.
Figure 6: Feature Trajectories. Trajectories of (random sampled) decoder vectors represent the directional change of features across training checkpoints. “Dark red” indicates features that are considered “formed,” i.e., they have gained semantic meaning and generally remain stable semantic meaning until the final state. “Blue” indicates features that are still unformed or in the initial stage. Conducted on Pythia-410m-deduped.
This continued geometric drift of semantically “formed” (Definition 5.2) features is vividly illustrated in Fig. 6: “formed” segments (dark red) still exhibit noticeable movement. This demonstrates that an individual feature’s direction can evolve even after its semantic identity has stabilized. It is indeed counter-intuitive that a single feature can maintain such clear semantic constancy while its geometric direction is still undergoing change, highlighting a distinct decoupling of semantic and directional stabilization at the individual feature level. This persistent directional adjustment of individual features reflects ongoing changing within the overall feature space. Such refinement of individual components is not contradictory to the system’s broader convergent phase; rather, it describes how individual features continue to optimize their geometric placement as the entire network settles towards a global equilibrium.
@@ 6 Conclusion
In this paper, we conduct a comprehensive mechanistic analysis of feature evolution in LLMs during training: (1) We propose SAE-Track, a novel method for efficiently obtaining a continual series of SAEs across training checkpoints, providing the foundation for detailed mechanistic studies. (2) We systematically characterize semantic evolution by identifying comprehensive patterns and phases. (3) We mechanistically investigate feature formation, modeling it as the geometric convergence of datapoints into localized regions. Our proposed progress measure captures this gradual process while effectively distinguishing the differing dynamics of token-level and concept-level features. (4) We analyze feature drift, finding directions undergo significant, three-phase adjustments. Counter-intuitively, this drift persists even after features are semantically formed, with full stabilization only occurring late in training. Our work provides a detailed understanding of how features evolve throughout training, demystifying training dynamics from the perspective of SAE-based analysis.
Appendix A Limitations
Model Architecture and Scale.
Our analysis is currently restricted to mid-scale autoregressive language models that offer complete and publicly available training checkpoints. Specifically, our work utilizes the Pythia suite [2] and the Stanford-CRFM GPT-2 small/medium models [16], both of which are natively supported by the TransformerLens library [25].
This focus is driven by significant constraints. Analyzing considerably larger models is computationally prohibitive due to the resource demands of processing numerous intermediate checkpoints. More critically, most contemporary state-of-the-art open source models lack comprehensive public releases of these intermediate training records. As detailed in Table 1, while some models, particularly those developed for scientific research, offer extensive checkpoints, a majority of prominent SOTA architectures provide limited or no such access. Current open-source practices often prioritize releasing final model weights but not the intermediate checkpoints vital for in-depth studies of pre-training dynamics and feature evolution.
Table 1: Public availability of intermediate training checkpoints. Only the first block (above the mid-rule) releases intermediate checkpoints. TransformerLens indicates native compatibility with the TransformerLens interpretability library. Only include intermediary checkpoints. Snapshots are only shared with selected researchers upon request; no public dump exists.
Model family Arch. / Lic. Size # ckpts TransformerLens
Pythia (deduped) [2] GPT-NeoX / Apache-2.0 70M–12B 154
CRFM GPT-2 [16] GPT-2 / Apache-2.0 124M / 355M 609
Amber-7B [22] LLaMA-7B / Apache-2.0 7B 360
Crystal-7B [22] LLaMA-7B / Apache-2.0 7B 143
BLOOM 176B [32] Transformer / RAIL 176B 21
OLMo [13] Transformer / Apache-2.0 1B / 7B $\sim$1454
OLMo2 [27] Transformer / Apache-2.0 7B / 13B $\sim$964
OPT 125M–175B [38] Transformer / NC 125M–175B request ✓(no intermediate)
LLaMA-3 [12] Transformer / Llama 3 8B–405B
Gemma 2 [34] Transformer / Gemma 2–27B
Qwen 3 [36] Transformer / Apache-2.0 0.6B–32B
Therefore, considering the limited availability of comprehensive SOTA checkpoints, the necessity of TransformerLens compatibility, and our computational capacity, the Pythia and Stanford-CRFM GPT-2 families were selected as the most suitable for this study.
Human Annotation and Semantic Interpretation.
In our study, the semantic labeling of features extracted by sparse autoencoders (SAEs) depends on the manual inspection of top-$k$ activating contexts, subsequently followed by the assignment of human-interpretable descriptors. Although this annotation method is standard within SAE literature[3, 35], it inherently introduces subjectivity and restricts scalability. Developing automated interpretation pipelines with LLM represents an important avenue for future research to address these limitations.
Appendix B Detailed SAE Formulation
This appendix provides the detailed mathematical formulation of the Sparse Autoencoders (SAEs) used in our work, following [3, 35].
Let $\mathbf{x}\in\mathbb{R}^{D}$ denote the input activations (e.g., from an LLM’s residual stream), where $D$ is the dimensionality of the activation space. The SAE learns a dictionary of $F$ features, and its operations can be described as follows:
Encoder.
Feature activations $f_{i}(\mathbf{x})$ for each feature $i$ are computed by the encoder:
Here, $\mathbf{W}^{\text{enc}}\in\mathbb{R}^{F\times D}$ are the encoder weights, $\mathbf{b}^{\text{enc}}\in\mathbb{R}^{F}$ are the encoder biases. The term $c\cdot\mathbf{b}^{\text{dec}}$ involves the decoder bias $\mathbf{b}^{\text{dec}}\in\mathbb{R}^{D}$ and a constant $c\in\{0,1\}$ that determines if the decoder bias is subtracted from the input before encoding.
Decoder.
The SAE reconstructs the input activations as $\widehat{\mathbf{x}}$ using the decoder:
where $\mathbf{W}^{\text{dec}}\in\mathbb{R}^{D\times F}$ are the decoder weights.
Loss Function.
The SAE is trained by minimizing a loss function $\mathcal{L}$ that typically combines a reconstruction error term and an L1 sparsity penalty on the feature activations:
where $\lambda$ is a hyperparameter controlling the strength of the sparsity penalty. The L1 penalty term, $\mathcal{L}_{1}$, can take different forms depending on constraints applied to the decoder weights:
$\mathcal{L}_{1}=\begin{cases}\sum_{i=1}^{F}|f_{i}(\mathbf{x})|&\text{(if decoder weights }\mathbf{W}^{\text{dec}}_{:,i}\text{ are unit-normalized)},\\ \sum_{i=1}^{F}|f_{i}(\mathbf{x})|\cdot\|\mathbf{W}^{\text{dec}}_{:,i}\|_{2}&\text{(if no unit norm constraint on decoder weights)}.\end{cases}$ (14)
In our experiments, we follow the setup where decoder weights are unit-normalized during training, thus using the first form of the L1 penalty.
Appendix C Detailed Derivation of the Training-Step Continuity Theorem